# Light Propagation in $\kappa$-Minkowski Space-Time: Gauge Ambiguities and Invariance

### v1

@@ 5 Finite gauge transformations
$\bar{\gamma}^{\mu}_{\sigma}(\Lambda)\partial_{\nu}\Lambda_{\mu}-\bar{\gamma}^{\mu}_{\nu}(\Lambda)\partial_{\sigma}\Lambda_{\mu}+\frac{1}{2}\bar{\gamma}^{\mu}_{\xi}(\Lambda)f^{\xi\chi}_{\omega}\,x^{\omega}\,\bar{\gamma}^{\zeta}_{\chi}(\Lambda)\,\partial_{\nu}\Lambda_{\mu}\,\partial_{\sigma}\Lambda_{\zeta}=0.$ (5.7)
As in the commutative case, any constant field satisfies this relation. Below, we shall work with the following one-parameter family of pure gauges:
@@ 7 Summary and concluding discussions
The gauge dependence of both the speed of light and the distances between charged particles is self-consistent in the following sense. Under the gauge transformations (5.12) and (6.5), which affect both fields and particle positions, the time taken by a signal to travel from one particle to another remains independent of the parameter $\alpha$ . By defining the unit of length in the standard way (6.6), we see that observers associated with gauge-related realizations of the dynamics will measure the same length and the same speed of light. The gauge independence of measurement results implies that gauges characterized by different values of $\alpha$ are physically equivalent.

### v2

@@ 5 Finite gauge transformations
$\bar{\gamma}^{\mu}_{\sigma}(\Lambda)\partial_{\nu}\Lambda_{\mu}-\bar{\gamma}^{\mu}_{\nu}(\Lambda)\partial_{\sigma}\Lambda_{\mu}+\frac{1}{2}\bar{\gamma}^{\mu}_{\xi}(\Lambda)f^{\xi\chi}_{\omega}\,x^{\omega}\,\bar{\gamma}^{\zeta}_{\chi}(\Lambda)\,\partial_{\nu}\Lambda_{\mu}\,\partial_{\sigma}\Lambda_{\zeta}=0,$ (5.7)
where $\bar{\gamma}$ stands for the inverse matrix of $\gamma$. As in the commutative case, any constant field satisfies this relation. Below, we shall work with the following one-parameter family of pure gauges:
@@ 7 Summary and concluding discussions
The gauge dependence of both the speed of light and the distances between charged particles is self-consistent in the following sense. Under the gauge transformations (5.12) and (6.5), which affect both fields and particle positions, the time taken by a signal to travel from one particle to another remains independent of the parameter $\beta$ . By defining the unit of length in the standard way (6.6), we see that observers associated with gauge-related realizations of the dynamics will measure the same length and the same speed of light. The gauge independence of measurement results implies that gauges characterized by different values of $\alpha$ are physically equivalent.
@@
Acknowledgements
The author is grateful to Vlad Kupriyanov, Fedele Lizzi, Alexey Sharapov, and Patrizia Vitale for valuable discussions on this and related papers.